OpenStudy (anonymous):

Simplify: (3x - 8) (2x - 2). show your work plz

OpenStudy (anonymous):

3x*2x=6x^2 3x*-2=6x -8*2x=-16x -8*-2=16 6x^2+6x-16x+16= 6x^2-10x+16

OpenStudy (anonymous):

96 i think